### v1

@@ 3.1 FRG formalism
We briefly lay out the basis of the FRG formalism, referring the reader to comprehensive introductions, such as [42, 43, 44, 45], for more details. The FRG is based on the idea of Wilsonian RG. Instead of averaging over the fluctuating fields $\Phi$ in the path integral (9) at all (momentum) scales simultaneously, one only integrates up to a certain RG scale $\kappa$. One then considers how this modified theory evolves under an infinitesimal change of the RG scale, which gives rise to a deterministic differential equation, where the role ot time is played by the RG scale, and the corresponding trajectories are in an abstract theory space, and connect a given microscopic model to its effective theory at large distance and time. From the technical point of view, this is achieved in FRG through the introduction of a scale-dependent ‘weight’ $\exp(-\Delta\mathcal{S}_{\kappa})$ in the functional integral to suppress fluctuations below the RG scale $\kappa$:
$\mathcal{Z}_{\kappa}=\int{\mathrm{D}}\Phi\exp\left(-\mathcal{S}[\Phi]-\Delta\mathcal{S}_{\kappa}[\Phi]+\int_{t,\mathbf{x}}\mathcal{J}\Phi\right),$ (20)
where $\Delta\mathcal{S}_{\kappa}[\Phi]\equiv\frac{1}{2}\int_{t,\mathbf{x},\mathbf{x}^{\prime}}\Phi(t,\mathbf{x})\mathcal{R}_{\kappa}(|\mathbf{x}-\mathbf{x}^{\prime}|)\Phi(t,\mathbf{x}^{\prime})$. The kernel $\mathcal{R}_{\kappa}$ is called ‘regulator’ and must satisfy the following properties (in Fourier space): (i) $\mathcal{R}_{\kappa}(\mathbf{p})\stackrel{{\scriptstyle\kappa\rightarrow\Lambda}}{{\sim}}\Lambda^{2}$ such that all fluctuations are frozen at the initial (microscopic) RG scale $\Lambda$; (ii) $\mathcal{R}_{\kappa}(\mathbf{p})\stackrel{{\scriptstyle\kappa\rightarrow 0}}{{\longrightarrow}}0$ such that all fluctuations are averaged over at the final (large-distance) RG scale; (iii) $\mathcal{R}_{\kappa}(\mathbf{p})\stackrel{{\scriptstyle|\mathbf{p}|\ll\kappa}}{{\sim}}\kappa^{2}$ such that at a scale $\kappa$, the contribution of the ‘slow’ modes is suppressed; (iv) $\mathcal{R}_{\kappa}(\mathbf{p})\stackrel{{\scriptstyle|\mathbf{p}|\gg\kappa}}{{\longrightarrow}}0$ such that at a scale $\kappa$, the ‘fast’ modes are unaltered and averaged over. In the presence of the regulator, the definition (12) is slightly modified to
@@
$\Gamma_{\kappa}[\Psi]+\Delta\mathcal{S}_{\kappa}[\Psi]=\underset{\mathcal{J}}{\sup}\left[-\mathcal{W}_{\kappa}[\mathcal{J}]+\int_{t,\mathbf{x}}\mathcal{J}\Psi\right]\,,$ (21)
@@
where ${\cal W}_{\kappa}=\ln{\cal Z}_{\kappa}$, and $\Gamma_{\kappa}$ is called the effective average action. In the large-$\kappa$ limit, the microscopic action $\Gamma_{\Lambda}\equiv\mathcal{S}$ is recovered since all the fluctuations are frozen, while in the small-$\kappa$ limit, the full effective action $\Gamma_{0}\equiv\Gamma$, encompassing all the fluctuations at all scales, is obtained [44, 45]. This provides $\Gamma_{\kappa}$ with a physical interpretation as the effective theory at scale $\kappa$, interpolating from the microscopic theory at $\kappa=\Lambda$ to the complete effective theory when $\kappa\to 0$. The very idea of the Wilsonian RG is thus realised through this formalism in a very intuitive way. The evolution of $\Gamma_{\kappa}$ with the RG scale in between these two limits is given by the exact Wetterich equation [46, 47, 48]
@@
where the trace means summation over all fields, $\Gamma_{\kappa}^{(2)}$ is the Hessian of $\Gamma$, and $\mathcal{G}_{\kappa}$ the propagator matrix.
The flow equation (22) is exact, but as it is a functional partial differential equation, it cannot be solved exactly in general. In the following, we present two complementary approaches to treat these equations for model A and investigate its inviscid limit. One of them consists in devising an Ansatz for $\Gamma_{\kappa}$, which allows one to project the flow onto a set of chosen renormalised functions, which are then numerically solved, as done in eg. Ref. [49]. This approach is particularly accurate to describe the small-momentum behaviour. It allows us to first qualitatively establish the fixed point structure of model A, evidencing the existence of the IFNS fixed point. The other approach consists in expanding the flow equation for the correlation function at large momentum. The leading order of this equation can then be closed exactly using the Ward identities stemming from the extended symmetries. Moreover, it can be solved analytically, yielding the exact asymptotic expression of the correlation function at large momentum. This approach is appropriate to describe the large-momentum behaviour. It allows us to show that $z=1$ exactly at the IFNS fixed point. Finally, to obtain a quantitative picture over the whole range of momenta, we consider an advanced numerical method which allows us to solve simultaneously the two sets (large- and small-momentum) of flow equations and compute the full correlation function and the associated half decay frequency $\omega_{1/2}(p)$. We start by presenting the large-momentum equations to establish the general result $z=1$, before turning to the small-momentum one to show the existence of the new IFNS fixed point.
@@ 3.2 Exact closure at large momentum
The exact FRG flow equations for any vertex function $\Gamma_{\kappa}^{(m,n)}$ can be obtained by taking functional derivatives of Eq. (22). It turns out that it can be simplified in a controlled way in the limit of large momenta. This originates in the very structure of these flow equations, in particular in the presence of the regulator term ${\partial}_{s}\mathcal{R}_{\kappa}(\mathbf{q})$ in the loop integral. It implies that the integral over the internal momentum $\mathbf{q}$ is cut to values $|\mathbf{q}|\lesssim\kappa$. Thus, in the limit where all the external momenta are large $|\mathbf{p}_{i}|\gg\kappa$ for all $i$, one has $|\mathbf{q}|\ll|\mathbf{p}_{i}|$. Because analyticity of the vertices is guaranteed at all scales $\kappa$ [45], the vertices appearing in the integrand can be Taylor expanded in powers of $\mathbf{q}$, since $|\mathbf{q}|$ is always compared to one of the $|\mathbf{p}_{i}|$ and $|\mathbf{q}|/|\mathbf{p}_{i}|\to 0$ for all $i$. This expansion becomes exact in the limit where all the $\mathbf{p}_{i}$ (and their partial sums) are infinite and in the vicinity of a fixed point [50, 51, 52].
@@
While the former property always holds, an exceptional feature occurs in the context of turbulence, which stems from the existence of the extended symmetries highlighted in Sec. 22.2. Indeed, the set of related Ward identities (15), (17) (and similar higher-order ones) appears to fix all the vertices with one zero momentum in terms of lower-order vertices. These are precisely the configurations which arise in the flow equations expanded at large momenta ($|\mathbf{q}|/|\mathbf{p}_{i}|\to 0$), which can thus be closed exactly using the Ward identities. Moreover, the resulting closed flow equations can be solved exactly in the vicinity of a fixed point. This exact closure was first achieved for the forced (non-equilibrium) Navier-Stokes equation in three, and then two dimensions, and for passively advected scalars [49, 33, 34, 53, 54]. It was also demonstrated for the one-dimensional, and $d$-dimensional, Burgers–KPZ equation [23, 55]. We refer the readers to these references or to the review [35] for the complete proof.
@@
The crucial point is that the exact FRG flow equations for the $n$-point correlation functions are very generic, as well as their large momentum expansion. Moreover, their exact closure only relies on the extended symmetries, and not on the details of the specific hydrodynamical equation, in particular not on the precise form of the noise or forcing term provided it has a Gaussian distribution, since this term does not enter the Ward identities (15), (17). In this respect, the large momentum equations of model A are the same as the ones obtained for the forced Navier-Stokes, except that there is only one independent two-point function because of the identity (19) stemming from time-reversal symmetry. We give it for the transverse correlation function $\bar{C}(\varpi,\mathbf{p})\equiv P_{\alpha\beta}^{\perp}(\mathbf{p})\bar{G}_{\alpha\beta}^{(2,0)}(\varpi,\mathbf{p})$ (the longitudinal part vanishes because of incompressibility) [34, 35]
@@
with $\partial_{s}\equiv\kappa\partial_{\kappa}$, and where ~$\tilde{\partial}_{s}$ only acts on the scale dependence of the regulator. This equation can be solved at a fixed point. Fourier transforming back to time delay, the solution reads [34, 35]
@@
where $\mu_{0}$ and $\mu_{\infty}$ are non-universal constants (related to the fixed-point values of $J_{\kappa}(0)$ and $\int_{\omega}J_{\kappa}(\omega)$ respectively) and $\tau_{c}$ is a characteristic time scale. This expression shows that the correlation function takes at small time delays a scaling form similar to (6) with the scaling variable $x=pt$, and thus $z=1$, different from the FNS value $z=2$. Hence, even in the viscous case where the small momenta (large distance) converge to the scaling form (6) with $z=2$, the large momenta exhibit a different scaling when the viscosity is small enough. Moreover, the solution (24) predicts a crossover at large time delays to yet another scaling regime, with an exponent $z=2$, but which differs from the viscous regime since $\mu_{\infty}$ does not depend on $\nu$, and it emerges at large momentum (not at large distances). Let us note that the solution (24) is an exact asymptotic result, but it does not give quantitative information about when (i.e. for how large momentum, and how small viscosity) it holds. This issue will be investigated in Sec. 33.4.
As previously mentioned, the result (24) is very general, it also holds for the large-momentum behaviour of the forced Navier-Stokes equation or passive scalar turbulence. It has been investigated in these contexts through direct numerical simulations. While the small-time $z=1$ decay is easily observed, eg. [28, 31], the large-time $z=2$ decay has remained elusive so far. Indeed, this regime turns out to be hard to resolve in numerical simulations since the initial Gaussian decay is very fast and the large-time correlations fall below the noise level. This regime has been evidenced so far only for a generalisation of Kraichnan’s model for passive scalars in Ref. [32]. We show in Sec. 4 that for model A, we manage to evidence both the small and the large time regimes in our direct numerical simulations.
@@ 3.3 FRG flow equations for model A
The asymptotic expression (24) holds for large $p$. However, the constants $\mu_{0}$, $\mu_{\infty}$ are related to the fixed-point values of $J_{\kappa}(\omega)$ in (23[34]. They hence primarily depend on the small $\mathbf{q}$ region, and their calculation requires the integration of the whole flow. An approximation which is accurate to describe the small momentum sector is to devise an ansatz for $\Gamma_{\kappa}$. Taking into account all the symmetries and associated non-renormalisation theorems, the appropriate ansatz reads [49]
@@
where in Fourier space $\bar{\nu}_{\kappa}(\omega,\mathbf{p})$ is a whole function of $\omega$ and $\mathbf{p}$, which identifies with the renormalised viscosity defined by (8). However, in contrast with perturbative calculations, the FRG allows one to compute the flow of this whole function (hence the F in FRG), and not only renormalised parameters as standard dynamical RG. The ansatz (25) encompasses the most general two-point functions, while it neglects all higher-order vertices but $\Gamma_{\kappa}^{(2,1)}$ which is present at the bare level and is not renormalised. It was shown to yield accurate results for 3D and 2D turbulence [35].
@@
×$\times\Big{[}\Big{(}2\mathbf{p}^{2}-\frac{(\mathbf{p}\cdot(\mathbf{p}+\mathbf{q}))^{2}}{(\mathbf{p}+\mathbf{q})^{2}}-\frac{(\mathbf{p}\cdot\mathbf{q})^{2}}{\mathbf{q}^{2}}\Big{)}(d-1)$
@@
where $R_{\kappa}$ denotes the component of the regulator matrix ${\cal R}_{\kappa}$, $R_{\kappa}\equiv[{\cal R}_{\kappa}]_{12}=[{\cal R}_{\kappa}]_{21}=-[{\cal R}_{\kappa}]_{22}/2$. with $\tilde{\nu}_{\kappa}(\varpi,\mathbf{q})=\mathbf{q}^{2}\bar{\nu}_{\kappa}(\varpi,\mathbf{q})+R_{\kappa}(\mathbf{q})$. The derivation of this equation can be found in Ref. [49], simply imposing at the end the constraints ensuing from time-reversal symmetry  [^1] The flow equation (26) corresponds to Eq. (B10) of Ref. [49], setting $f_{\perp}^{D}(p)=\chi f_{\perp}^{\nu}(\mathbf{p})$ and $N_{\kappa}=\chi R_{\kappa}$ to preserve time-reversal symmetry, and restoring the frequency arguments $f_{\perp}^{\nu}(\mathbf{q})\to\tilde{\nu}_{\kappa}(\varpi,\mathbf{q})$, $f_{\perp}^{\nu}(\mathbf{p}+\mathbf{q})\to\tilde{\nu}_{\kappa}(\omega+\varpi,\mathbf{p}+\mathbf{q})$. Note that a sign typo in the last line of (B10) has been corrected.. Note that a further simplification is used for the numerical resolution of this equation, which consists in neglecting the frequency dependence of $\tilde{\nu}_{\kappa}$ within the integral, which amounts to the replacement $\tilde{\nu}(\varpi,\mathbf{q})\simeq\tilde{\nu}_{\kappa}(0,\mathbf{q})$ in the integrand.

### v2

@@ 3.1 FRG formalism
We briefly lay out the basis of the FRG formalism, referring the reader to comprehensive introductions, such as [47, 48, 49, 50], for more details. The FRG is based on the idea of Wilsonian RG. Instead of averaging over the fluctuating fields $\Phi$ in the path integral (9) at all (momentum) scales simultaneously, one only integrates them down to a certain RG scale $\kappa$. One then considers how this modified theory evolves under an infinitesimal change of the RG scale, which gives rise to a deterministic differential equation, akin a dynamical system, where the role ot time is played by the RG scale, and the physical trajectories are replaced by trajectories in an abstract theory space. Such a trajectory connects a given microscopic model to its effective theory at large distance and time. In analogy with a fixed point in the dynamical flow which corresponds to a stationary (time-independent) state, a fixed point in the RG flow corresponds to a scale-invariant state.
From the technical point of view, the progressive averaging of fluctuations is achieved in FRG through the introduction of a scale-dependent ‘weight’ $\exp(-\Delta\mathcal{S}_{\kappa})$ in the functional integral to suppress fluctuations below the RG scale $\kappa$:
$\mathcal{Z}_{\kappa}=\int{\mathrm{D}}\Phi\exp\left(-\mathcal{S}[\Phi]-\Delta\mathcal{S}_{\kappa}[\Phi]+\int_{t,\mathbf{x}}\mathcal{J}_{i}\Phi_{i}\right),$ (20)
where $\Delta\mathcal{S}_{\kappa}[\Phi]\equiv\frac{1}{2}\int_{t,\mathbf{x},\mathbf{x}^{\prime}}\Phi_{i}(t,\mathbf{x})\mathcal{R}_{\kappa,ij}(|\mathbf{x}-\mathbf{x}^{\prime}|)\Phi_{j}(t,\mathbf{x}^{\prime})$. The kernel $\mathcal{R}_{\kappa}$ is called ‘regulator’. Its precise form is unimportant, as long as all its elements satisfy the following set of properties (in Fourier space):
(i) $\mathcal{R}_{\kappa,ij}(\mathbf{p})\stackrel{{\scriptstyle\kappa\rightarrow\Lambda}}{{\sim}}\Lambda^{2}$ such that all fluctuations are frozen at the initial (microscopic) RG scale $\Lambda$;
(ii) $\mathcal{R}_{\kappa,ij}(\mathbf{p})\stackrel{{\scriptstyle\kappa\rightarrow 0}}{{\longrightarrow}}0$ such that all fluctuations are averaged over at the final (macroscopic) RG scale;
(iii) $\mathcal{R}_{\kappa,ij}(\mathbf{p})\stackrel{{\scriptstyle|\mathbf{p}|\ll\kappa}}{{\sim}}\kappa^{2}$ such that at a scale $\kappa$, the contribution of the ‘slow’ modes is suppressed;
(iv) $\mathcal{R}_{\kappa,ij}(\mathbf{p})\stackrel{{\scriptstyle|\mathbf{p}|\gg\kappa}}{{\longrightarrow}}0$ such that at a scale $\kappa$, the ‘fast’ modes are unaltered and averaged over,
where we recall that $\Lambda^{-1}$ denotes the microscopic length scale at which the hydrodynamical model is defined. In the presence of the regulator, the Legendre transform (12) is slightly modified to
@@
$\Gamma_{\kappa}[\Psi]+\Delta\mathcal{S}_{\kappa}[\Psi]=\underset{\mathcal{J}}{\sup}\left[-\mathcal{W}_{\kappa}[\mathcal{J}]+\int_{t,\mathbf{x}}\mathcal{J}_{i}\Psi_{i}\right]\,$, (21)
@@
where ${\cal W}_{\kappa}=\ln{\cal Z}_{\kappa}$, and $\Gamma_{\kappa}$ is called the effective average action. In the large-$\kappa$ limit, $\Gamma_{\kappa}$ identifies with the microscopic action $\Gamma_{\Lambda}\equiv\mathcal{S}$ since all fluctuations are frozen, while in the small-$\kappa$ limit, the full effective action $\Gamma_{0}\equiv\Gamma$, encompassing all fluctuations at all scales, is obtained [49, 50]. This provides $\Gamma_{\kappa}$ with a physical interpretation as the effective theory at scale $\kappa$, interpolating from the microscopic theory at $\kappa=\Lambda$ to the complete effective theory when $\kappa\to 0$. The very idea of the Wilsonian RG is thus realised through this formalism in a consistent and efficient way. The evolution of $\Gamma_{\kappa}$ with the RG scale in between these two limits is given by the exact Wetterich equation [51, 52, 53]
@@
where the trace means summation over all fields, $\Gamma_{\kappa}^{(2)}$ is the Hessian of $\Gamma$, and $\mathcal{G}_{\kappa}$ the full (functional) propagator of the theory.
The flow equation (22) is exact, but as it is a functional partial differential equation, it cannot be solved exactly in general. Several approximation schemes are available to study it, which are non-perturbative and can be controlled and improved in a systematic way [50]. In the following, we present two complementary approaches to treat these equations for model A and investigate its inviscid limit. One of them consists in devising an Ansatz for $\Gamma_{\kappa}$, which allows one to project the flow onto a set of chosen renormalised functions, which are then numerically solved, as done in eg. Ref. [54]. This approach is particularly accurate to describe the small-momentum behaviour. It allows us to first qualitatively establish the fixed point structure of model A, evidencing the existence of the IFNS fixed point. The other approach consists in expanding the flow equation for the velocity correlation function at large momentum. It turns out that the leading order of this expansion can be closed exactly using the Ward identities stemming from the extended symmetries. Moreover, it can be solved analytically, yielding the exact asymptotic expression of the correlation function at large momentum. This approach is thus appropriate to describe the large-momentum behaviour. It allows us to show in particular that $z=1$ exactly at the IFNS fixed point. Finally, to obtain a quantitative picture over the whole range of momenta, we consider an advanced numerical method which allows us to solve simulaneously the two sets (large- and small-momentum) of flow equations and compute the full correlation function and the associated half decay frequency $\omega_{1/2}(p)$. We start by presenting the large-momentum equations to establish the general result $z=1$, before turning to the small-momentum one to show the existence of the new IFNS fixed point.
@@ 3.2 Exact closure at large momentum
The exact FRG flow equations for any vertex function $\Gamma_{\kappa}^{(m,n)}$ can be obtained by taking $m$ functional derivatives of Eq. (22) with respect to $\mathbf{u}$ fields and $n$ with respect to $\bar{\mathbf{u}}$ fields. It turns out that they can be simplified in a controlled way in the limit of large momenta. This originates in the very structure of these flow equations, in particular in the presence of the regulator term ${\partial}_{s}\mathcal{R}_{\kappa}(\mathbf{q})$ in the loop integral. It implies that the integral over the internal momentum $\mathbf{q}$ is cut to values $|\mathbf{q}|\lesssim\kappa$. Thus, in the limit where all the external momenta are large $|\mathbf{p}_{i}|\gg\kappa$ for all $i=1,\dots,m+n$, one has $|\mathbf{q}|\ll|\mathbf{p}_{i}|$. Because analyticity of the vertices is guaranteed at all scales $\kappa$ by construction of the FRG formalism [50], the vertices appearing in the flow equation can be Taylor expanded in powers of $\mathbf{q}$, since $|\mathbf{q}|$ is always compared to one of the $|\mathbf{p}_{i}|$ and $|\mathbf{q}|/|\mathbf{p}_{i}|\to 0$ for all $i$. This expansion becomes exact in the limit where all the $\mathbf{p}_{i}$ (and their partial sums) tend to infinity and in the vicinity of a fixed point [55, 56, 57]. In practice, for turbulence, the RG flow essentially stops when $\kappa\lesssim L^{-1}$ where $L$ is the integral scale, such that the large momentum condition becomes $|\mathbf{p}_{i}|\gg L^{-1}$.
@@
While the exactness of the large momentum expansion of the FRG flow always holds, an exceptional feature occurs in the context of turbulence, which stems from the existence of the extended symmetries highlighted in Sec. 22.2. Indeed, the set of related Ward identities (15), (17) (and the higher-order ones) appears to fix all the vertices with one zero momentum (associated with a velocity or a response velocity) in terms of lower-order vertices. These are precisely the configurations which arise in the flow equations expanded at large momenta ($|\mathbf{q}|/|\mathbf{p}_{i}|\to 0$), which can thus be closed exactly using the Ward identities. Moreover, the resulting closed flow equations can be solved analytically in the vicinity of a fixed point. This exact closure was first achieved for the forced (non-equilibrium) Navier-Stokes equation in three, and then two dimensions, and for passively advected scalars [54, 37, 38, 58, 59]. It was also demonstrated for the one-dimensional, and $d$-dimensional, Burgers–KPZ equation [28, 60]. We refer the readers to these references or to the review [39] for the complete proof.
@@
The crucial point is that the Wetterich FRG flow equations for the $n$-point correlation functions are very generic, as well as their large momentum expansion. Moreover, their exact closure only relies on the extended symmetries, and not on the details of the specific hydrodynamical equation, in particular not on the precise form of the noise or forcing term, provided it has a Gaussian distribution, since this term does not enter the Ward identities (15), (17). In this respect, the large momentum equations of model A are the same as the ones obtained for the forced Navier-Stokes, except that there is only one independent two-point function because of the identity (19) stemming from time-reversal symmetry. We give it for the transverse correlation function $\bar{C}(\varpi,\mathbf{p})\equiv P_{\alpha\beta}^{\perp}(\mathbf{p})\bar{G}_{\alpha\beta}^{(2,0)}(\varpi,\mathbf{p})$ (the longitudinal part vanishes because of incompressibility) [38, 39]
@@
with $\partial_{s}\equiv\kappa\partial_{\kappa}$, and where $\tilde{\partial}_{s}$ only acts on the scale dependence of the regulator. This equation can be solved at a fixed point. Fourier transforming back to time delay, the solution reads [38, 39]
@@
where $\mu_{0}$ and $\mu_{\infty}$ are non-universal constants (related to the fixed-point values of $J_{\kappa}(0)$ and $\int_{\omega}J_{\kappa}(\omega)$ respectively) and $\tau_{c}$ is a characteristic time scale. This expression shows that the correlation function takes at small time delays a scaling form similar to (6) with the scaling variable $x=pt$, and thus $z=1$, different from the FNS value $z=2$. Hence, even in the viscous case where the small momenta (large distance) converge to the scaling form (6) with $z=2$, the large momenta exhibit a different scaling when the viscosity is small enough. Moreover, the solution (24) predicts a crossover at large time delays to yet another scaling regime, with an exponent $z=2$. However, this regime differs from the viscous diffusive regime since $\mu_{\infty}$ does not depend on $\nu$, the scaling form is different from (7), and it emerges at large momentum (not at large distance). Let us note that the solution (24) is an exact asymptotic result, but it does not give quantitative information about when (i.e. for how large momentum, and how small viscosity) it holds. Indeed, for actual numbers, one would have to integrate the whole flow (not only its asymptotic form at large $p$) to calculate $J_{\kappa}(\omega)$ for the precise model studied. This quantity is in particular sensitive to the form of the noise or forcing. This issue will be investigated in Sec. 33.4.
As previously mentioned, the result (24) is very general, it also holds for the large-momentum behaviour of the forced Navier-Stokes equation or passive scalar turbulence. It has been investigated in these contexts through direct numerical simulations. While the small-time $z=1$ decay is easily observed, eg. [33, 36], the large-time $z=2$ decay has remained elusive so far. Indeed, this regime turns out to be hard to resolve in numerical simulations since the initial Gaussian decay is very fast and the large-time correlations fall below the noise level. This regime has been evidenced so far only for a generalisation of Kraichnan’s model for passive scalars in Ref. [61]. We show in Sec. 4 that for model A, we manage to evidence both the small and the large time regimes in our direct numerical simulations, which is therefore an important result of the present work.
@@ 3.3 FRG flow equations for model A
The asymptotic expression (24) holds for large $p$. However, as previously mentioned, the constants $\mu_{0}$, $\mu_{\infty}$ are related to the fixed-point values of $J_{\kappa}(\omega)$ in (23[38]. They hence primarily depend on the small $p$ region, and their calculation requires the integration of the whole flow. An approximation which is accurate to describe the small momentum sector is to devise an ansatz for $\Gamma_{\kappa}$. Taking into account all the symmetries and associated non-renormalisation theorems, the appropriate ansatz reads [54]
@@
where in Fourier space $\bar{\nu}_{\kappa}(\omega,\mathbf{p})$ is a whole function of $\omega$ and $\mathbf{p}$, which identifies with the renormalised viscosity defined by (8). Note that the same function $\bar{\nu}_{\kappa}$ appears for both the dissipation and the noise term, as a consequence of the time-reversal symmetry (19). However, in contrast with perturbative calculations, the FRG allows one to compute the flow of this whole function (hence the F in FRG), and not only renormalised parameters as standard dynamical RG. The ansatz (25) encompasses the most general two-point functions, while it neglects all higher-order vertices but $\Gamma_{\kappa}^{(2,1)}$ which is present at the bare level and is not renormalised. It was shown to yield accurate results for 3D and 2D turbulence [39].
@@
$\times\Big[\Big(2\mathbf{p}^{2}-\frac{(\mathbf{p}\cdot(\mathbf{p}+\mathbf{q}))^{2}}{(\mathbf{p}+\mathbf{q})^{2}}-\frac{(\mathbf{p}\cdot\mathbf{q})^{2}}{\mathbf{q}^{2}}\Big)(d-1)$
@@
where $R_{\kappa}$ denotes the component of the regulator matrix ${\cal R}_{\kappa}$ defined as $R_{\kappa}\equiv{\cal R}_{\kappa,12}={\cal R}_{\kappa,21}=-{\cal R}_{\kappa,22}/2$, and where $\tilde{\nu}_{\kappa}(\varpi,\mathbf{q})=\mathbf{q}^{2}\bar{\nu}_{\kappa}(\varpi,\mathbf{q})+R_{\kappa}(\mathbf{q})$. The derivation of this equation can be found in Ref. [54], simply imposing at the end the constraints ensuing from time-reversal symmetry  [^2] The flow equation (26) corresponds to Eq. (B10) of Ref. [54], setting $f_{\perp}^{D}(p)=\chi f_{\perp}^{\nu}(\mathbf{p})$ and $N_{\kappa}=\chi R_{\kappa}$ to preserve time-reversal symmetry, and restoring the frequency arguments $f_{\perp}^{\nu}(\mathbf{q})\to\tilde{\nu}_{\kappa}(\varpi,\mathbf{q})$, $f_{\perp}^{\nu}(\mathbf{p}+\mathbf{q})\to\tilde{\nu}_{\kappa}(\omega+\varpi,\mathbf{p}+\mathbf{q})$. Note that a sign typo in the last line of (B10) has been corrected.. Note that a further simplification is used for the numerical resolution of this equation, which consists in neglecting the frequency dependence of $\tilde{\nu}_{\kappa}$ within the integral, which amounts to the replacement $\tilde{\nu}(\varpi,\mathbf{q})\simeq\tilde{\nu}_{\kappa}(0,\mathbf{q})$ in the integrand.